Exercise 1-6 (Algo) Traditional and Contribution Format Income Statements [LO1-6] Cherokee Inc. is a merchandiser that provided the following information: Amount Number of units sold 12,000 Selling price per unit $ 16 Variable selling expense per unit $ 1 Variable administrative expense per unit $ 2 Total fixed selling expense $ 18,000 Total fixed administrative expense $ 14,000 Beginning merchandise inventory $ 12,000 Ending merchandise inventory $ 23,000 Merchandise purchases $ 87,000 Required: 1. Prepare a traditional income statement. 2. Prepare a contribution format income statement.
Answer #1

Answer 1 and 2:

Sales Revenue = 12000 * 16 = $192,000

Variable selling expense = 12000 * $1 = $12,000

Variable administrative expense = 12000 * $2 = $24,000

Cost of goods sold = Beginning merchandise inventory + Merchandise purchases - Ending merchandise inventory

= 12000 + 87000 - 23000

= $76,000

Selling expenses = 12000 + 18000 = $30,000

Administrative expense = 24000 + 14000 = $38,000

Cherokee, Inc. Traditional Income Statement $192,000 $76,000 $116,000 Sales Revenue Cost of goods sold Gross margin Selling a